OpenStudy (anonymous):

What is the general form of the equation of a circle with a center at the (4, 2) and a radius of 9?

OpenStudy (jdoe0001):

so, what's the equation for a circle, from your formulas sheet?

OpenStudy (anonymous):

(x -h)^2 + (y - k)^2 =r^2?

OpenStudy (austinl):

The general equation for a circle is \[(x-h)^2+(y-k)^2=r^2\] h is the transformation going on the x axis, and k is the transformation on the y axis.

OpenStudy (jdoe0001):

well, then just PLUG IN the values :), what you were given is (h,k)

OpenStudy (jdoe0001):

and "r" too, so, you have everything :)

OpenStudy (anonymous):

oooh lol makes sense thank you :)

OpenStudy (austinl):

So, to move a circle to the right, or positive x direction h must be negative. And to move a circle up, or positive y direction, k must be negative.

OpenStudy (anonymous):

oh ok I get it now ty :)

OpenStudy (austinl):

So we are moving the circle right 4 and up 2, both positive. So, what does that make our formula?

OpenStudy (anonymous):

is it (x – 4)^2 + (y – 2)^2 = 81?

OpenStudy (austinl):

It is indeed! Good Job!

OpenStudy (anonymous):

awesome thanks so much :)

OpenStudy (austinl):

No problem!